[Hallc_running] [New Logentry] RC Summary July 9th

billlee at jlab.org billlee at jlab.org
Wed Jul 9 13:00:02 EDT 2025


Logentry Text:
--
Summary of Recent Activities:

Since the last RC update, we have successfully recovered from the BSY dipole trip, which resulted in approximately 3.5 hours of lost beam time. MCC plans to replace the relevent breaker to improve long-term stability. Data collection resumed and continued in production mode. This morning ~5 am, there was a LAD trip incident—multiple paddles required resetting (https://logbooks.jlab.org/entry/4414486). Beam was switched as scheduled at 6:45 a.m., and Hall C entered restricted access at approximately 8:15 a.m. Magnet ramp-down was completed by 8:30 a.m.

Maintenance Day Status (as of noon):

The following tasks have been completed:
1. LAD GEM stack survey work (by Survey Group, supported by Hall C Tech and GEM experts)
2. 10:00 a.m. tour (by Sanghwa)
3. 11:00 a.m. walkthrough (by Sam)
4. HMS and SHMS hodoscope diagnostics (by Simona & Bill Henry)
5. Inspection of the HMS JT valve and rotation control (by the spectrometer group)
6. SHMS HV control laptop surveyed and sent to IT to resolve login issues. (Requested by Simona and Bill)

Remaining tasks for the afternoon:
1. Remove the hand guard rail (Hall Tech)
2. Unlock power supplies for all components (Ellen)
3. Validate the DAQ (by GEM experts)
4. Ramp magnets (by shift crew).
5. Magnet checks to be performed (by Heidi)

Hall C remains on schedule to close up by 4:00 p.m.

RC Meetings:
Our daily RC meeting will take place today at 2:30 p.m. Connection details are available here: https://logbooks.jlab.org/entry/4346862

---

This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4414716
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20250709/87754245/attachment.htm>


More information about the Hallc_running mailing list